# Heavy Mesons to Charmed Tetraquark Decays

### v1

@@ V.2.1 $\overline{B}\to TP$ decay amplitudes
$3^{*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}\pi^{0}$ $-\frac{1}{\sqrt{2}}(C_{TP}-C_{PT,1})$ $3^{*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}\eta$ $-\frac{c_{\phi^{\prime}}}{\sqrt{2}}(C_{TP}+C_{PT,1})-s_{\phi^{\prime}}C_{T}$
$3^{*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}\eta^{\prime}$ $-\frac{s_{\phi^{\prime}}}{\sqrt{2}}(C_{TP}+C_{PT,1})+c_{\phi^{\prime}}C_{T}$ $4^{*}$ $\overline{B}{}^{0}\to T^{0}_{cs}K^{0}$ $-C_{PT,1}+E_{\bar{B}T}$
$5$ $\overline{B}{}^{0}\to T^{+}_{c}\pi^{-}$ $-\frac{1}{\sqrt{2}}(C_{PT,1}-E_{\bar{B}T})$ $6$ $B^{-}\to T^{0}_{c}\pi^{-}$ $\frac{1}{\sqrt{2}}(C_{T}+C_{TP})$
$6$ $\overline{B}{}^{0}\to T^{0}_{c}\pi^{0}$ $\frac{1}{2}(-C_{T}-C_{TP}+C_{PT,1}-E_{\bar{B}T})$ $6$ $\overline{B}{}^{0}\to T^{0}_{c}\eta$ $\frac{c_{\phi^{\prime}}}{2}(C_{T}-C_{TP}-C_{PT,1}+E_{\bar{B}T})+\frac{s_{\phi^{\prime}}}{\sqrt{2}}E_{\bar{B}T}$
$6$ $\overline{B}{}_{s}^{0}\to T^{0}_{c}K^{0}$ $\frac{1}{\sqrt{2}}(C_{T}+C_{PT,1})$ $6$ $\overline{B}{}^{0}\to T^{0}_{c}\eta^{\prime}$ $\frac{s_{\phi^{\prime}}}{2}(C_{T}-C_{TP}-C_{PT,1}+E_{\bar{B}T})-\frac{c_{\phi^{\prime}}}{\sqrt{2}}E_{\bar{B}T}$
$7$ $\overline{B}{}^{0}\to T^{\prime\prime+}_{c}\pi^{-}$ $-\frac{1}{\sqrt{2}}(2T_{P}+C_{PT,1}-E_{\bar{B}T}+2E_{\bar{B}P})$ $8$ $B^{-}\to T^{\prime\prime 0}_{c}\pi^{-}$ $-\frac{1}{\sqrt{2}}(2T_{P}+C_{T}+C_{TP}+2C_{PT,2})$
$8$ $\overline{B}{}^{0}\to T^{\prime\prime 0}_{c}\eta$ $\frac{c_{\phi^{\prime}}}{2}(-C_{T}+C_{TP}+C_{PT,1}-2C_{PT,2}$ $8$ $\overline{B}{}^{0}\to T^{\prime\prime 0}_{c}\eta^{\prime}$ $\frac{s_{\phi^{\prime}}}{2}(-C_{T}+C_{TP}+C_{PT,1}-2C_{PT,2}$
$-E_{\bar{B}T}-2E_{\bar{B}P})+\frac{s_{\phi^{\prime}}}{\sqrt{2}}E_{\bar{B}T}$ $-E_{\bar{B}T}-2E_{\bar{B}P})-\frac{c_{\phi^{\prime}}}{\sqrt{2}}E_{\bar{B}T}$
$8$ $\overline{B}{}^{0}\to T^{\prime\prime 0}_{c}\pi^{0}$ $\frac{1}{2}(C_{T}+C_{TP}-C_{PT,1}+2C_{PT,2}$ $8$ $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}K^{0}$ $-\frac{1}{\sqrt{2}}(C_{T}-C_{PT,1}+2C_{PT,2})$
$+E_{\bar{B}T}-2E_{\bar{B}P})$
$9$ $\overline{B}{}^{0}\to T^{\prime\prime+}_{c\bar{s}}K^{-}$ $\frac{1}{\sqrt{2}}(C_{TP}+E_{\bar{B}T}-2E_{\bar{B}P})$ $9$ $\overline{B}{}_{s}^{0}\to T^{\prime\prime+}_{c\bar{s}}\pi^{-}$ $-\frac{1}{\sqrt{2}}(2T_{P}+C_{TP}+C_{PT,1})$
# Mode $A^{\prime}(\overline{B}_{q}\to T_{cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}]}P)$ Mode $A^{\prime}(\overline{B}_{q}\to T_{cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}]}P)$
$2^{*}$ $\overline{B}{}_{s}^{0}\to T^{+}_{c\bar{s}}K^{-}$ $-\frac{1}{\sqrt{2}}(C^{\prime}_{PT,1}-E^{\prime}_{\bar{B}T})$ $3^{*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}\overline{K}^{0}$ $-C^{\prime}_{PT,1}+E^{\prime}_{\bar{B}T}$
$4^{*}$ $B^{-}\to T^{0}_{cs}\pi^{-}$ $C^{\prime}_{T}+C^{\prime}_{TP}$ $4^{*}$ $\overline{B}{}^{0}\to T^{0}_{cs}\pi^{0}$ $-\frac{1}{\sqrt{2}}(C^{\prime}_{T}+C^{\prime}_{TP})$
$4^{*}$ $\overline{B}{}^{0}\to T^{0}_{cs}\eta$ $\frac{c_{\phi^{\prime}}}{\sqrt{2}}(C^{\prime}_{T}-C^{\prime}_{TP})+s_{\phi^{\prime}}C^{\prime}_{PT,1}$ $4^{*}$ $\overline{B}{}^{0}\to T^{0}_{cs}\eta^{\prime}$ $\frac{s_{\phi^{\prime}}}{\sqrt{2}}(C^{\prime}_{T}-C^{\prime}_{TP})-c_{\phi^{\prime}}C^{\prime}_{PT,1}$
$4^{*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{cs}K^{0}$ $C^{\prime}_{T}+E^{\prime}_{\bar{B}T}$ $5$ $\overline{B}{}^{0}\to T^{+}_{c}K^{-}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{TP}-C^{\prime}_{PT,1})$
$5$ $\overline{B}{}_{s}^{0}\to T^{+}_{c}\pi^{-}$ $-\frac{1}{\sqrt{2}}(C^{\prime}_{TP}-E^{\prime}_{\bar{B}T})$ $6$ $B^{-}\to T^{0}_{c}K^{-}$ $-\frac{1}{\sqrt{2}}(C^{\prime}_{T}+C^{\prime}_{TP})$
$6$ $\overline{B}{}^{0}\to T^{0}_{c}\overline{K}^{0}$ $-\frac{1}{\sqrt{2}}(C^{\prime}_{T}+C^{\prime}_{PT,1})$ $6$ $\overline{B}{}_{s}^{0}\to T^{0}_{c}\pi^{0}$ $\frac{1}{2}(C^{\prime}_{TP}-E^{\prime}_{\bar{B}T})$
$6$ $\overline{B}{}_{s}^{0}\to T^{0}_{c}\eta$ $\frac{c_{\phi^{\prime}}}{2}(C^{\prime}_{TP}+E^{\prime}_{\bar{B}T})$ $6$ $\overline{B}{}_{s}^{0}\to T^{0}_{c}\eta^{\prime}$ $\frac{s_{\phi^{\prime}}}{2}(C^{\prime}_{TP}+E^{\prime}_{\bar{B}T})$
$+\frac{s_{\phi^{\prime}}}{\sqrt{2}}(C^{\prime}_{T}-C^{\prime}_{PT,1}+E^{\prime}_{\bar{B}T})$ $-\frac{c_{\phi^{\prime}}}{\sqrt{2}}(C^{\prime}_{T}-C^{\prime}_{PT,1}+E^{\prime}_{\bar{B}T})$
$7$ $\overline{B}{}^{0}\to T^{\prime\prime+}_{c}K^{-}$ $-\frac{1}{\sqrt{2}}(2T^{\prime}_{P}+C^{\prime}_{TP}+C^{\prime}_{PT,1})$ $7$ $\overline{B}_{s}{}^{0}\to T^{\prime\prime+}_{c}\pi^{-}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{TP}+E^{\prime}_{\bar{B}T}-2E^{\prime}_{\bar{B}P})$
$8$ $B^{-}\to T^{\prime\prime 0}_{c}K^{-}$ $-\frac{1}{\sqrt{2}}(2T^{\prime}_{P}+C^{\prime}_{T}+C^{\prime}_{TP}+2C^{\prime}_{PT,2})$ $8$ $\overline{B}{}^{0}\to T^{\prime\prime 0}_{c}\overline{K}^{0}$ $-\frac{1}{\sqrt{2}}(C^{\prime}_{T}-C^{\prime}_{PT,1}+2C^{\prime}_{PT,2})$
$8$ $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}\eta$ $\frac{c_{\phi^{\prime}}}{2}(C^{\prime}_{T}-E^{\prime}_{\bar{B}T}-2E^{\prime}_{\bar{B}P})$ $8$ $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}\eta^{\prime}$ $\frac{s_{\phi^{\prime}}}{2}(C^{\prime}_{T}-E^{\prime}_{\bar{B}T}-2E^{\prime}_{\bar{B}P})$
$+\frac{s_{\phi^{\prime}}}{\sqrt{2}}(C^{\prime}_{T}-C^{\prime}_{PT,1}+2C^{\prime}_{PT,2}+E^{\prime}_{\bar{B}T})$ $-\frac{c_{\phi^{\prime}}}{\sqrt{2}}(C^{\prime}_{T}-C^{\prime}_{PT,1}+2C^{\prime}_{PT,2}+E^{\prime}_{\bar{B}T})$
$8$ $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}\pi^{0}$ $\frac{1}{2}(C^{\prime}_{TP}+E^{\prime}_{\bar{B}T}-2E^{\prime}_{\bar{B}P})$ $9$ $\overline{B}{}_{s}^{0}\to T^{\prime\prime+}_{c\bar{s}}K^{-}$ $-\frac{1}{\sqrt{2}}(2T^{\prime}_{P}+C^{\prime}_{PT,1}-E^{\prime}_{\bar{B}T}+2E^{\prime}_{\bar{B}P})$
@@
$2^{\prime*}$ $\overline{B}{}^{0}\to T^{+}_{c\bar{s}}K^{-}$ $-\frac{1}{\sqrt{2}}(C_{TP}-E_{\bar{B}T})$
$2^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{+}_{c\bar{s}}\pi^{-}$ $-\frac{1}{\sqrt{2}}(C_{TP}-C_{PT,1})$
@@
$3^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}\eta$ $-s_{\phi^{\prime}}C_{T}+\frac{c_{\phi^{\prime}}}{\sqrt{2}}(C_{PT,1}+C_{TP})$
$3^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}\eta^{\prime}$ $c_{\phi^{\prime}}C_{T}+\frac{s_{\phi^{\prime}}}{\sqrt{2}}(C_{PT,1}+C_{TP})$
@@
$11^{\prime*}$ $B^{-}\to T^{0}_{c}\pi^{-}$ $-\sqrt{\frac{2}{3}}(C_{T}+C_{TP}-C_{PT,1})$
@@
$11^{\prime*}$ $\overline{B}{}^{0}\to T^{0}_{c}\eta$ $-\frac{c_{\phi^{\prime}}}{\sqrt{3}}(C_{T}+C_{TP}+C_{PT,1})$
$11^{\prime*}$ $\overline{B}{}^{0}\to T^{0}_{c}\eta^{\prime}$ $-\frac{s_{\phi^{\prime}}}{\sqrt{3}}(C_{T}+C_{TP}+C_{PT,1})$
$11^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c}K^{0}$ $-\sqrt{\frac{2}{3}}C_{T}$
@@
$12^{\prime*}$ $B^{-}\to T^{-}_{c}\eta$ $c_{\phi^{\prime}}(C_{T}+C_{TP}+C_{PT,1})$
$12^{\prime*}$ $B^{-}\to T^{-}_{c}\eta^{\prime}$ $s_{\phi^{\prime}}(C_{T}+C_{TP}+C_{PT,1})$
@@
$14^{\prime}$ $\overline{B}{}^{0}\to T^{\prime 0}_{c}\pi^{0}$ $-\frac{1}{2\sqrt{6}}(C_{T}-C_{TP}+C_{PT,1}-E_{\bar{B}T})$
$14^{\prime}$ $\overline{B}{}^{0}\to T^{\prime 0}_{c}\eta$ $\frac{c_{\phi^{\prime}}}{2\sqrt{6}}(C_{T}+C_{TP}+C_{PT,1}+3E_{\bar{B}T})+\frac{\sqrt{3}s_{\phi^{\prime}}}{2}E_{\bar{B}T}$
$14^{\prime}$ $\overline{B}{}^{0}\to T^{\prime 0}_{c}\eta^{\prime}$ $\frac{s_{\phi^{\prime}}}{2\sqrt{6}}(C_{T}+C_{TP}+C_{PT,1}+3E_{\bar{B}T})-\frac{\sqrt{3}c_{\phi^{\prime}}}{2}E_{\bar{B}T}$
@@
$17^{\prime}$ $\overline{B}{}^{0}\to T^{\prime\prime 0}_{c}\eta$ $\frac{c_{\phi^{\prime}}}{2\sqrt{2}}(C_{T}+C_{TP}+C_{PT,1}+4C_{PT,2}+3E_{\bar{B}T}+4E_{\bar{B}P})-\frac{s_{\phi^{\prime}}}{2}E_{\bar{B}T}$
$17^{\prime}$ $\overline{B}{}^{0}\to T^{\prime\prime 0}_{c}\eta^{\prime}$ $\frac{s_{\phi^{\prime}}}{2\sqrt{2}}(C_{T}+C_{TP}+C_{PT,1}+4C_{PT,2}+3E_{\bar{B}T}+4E_{\bar{B}P})+\frac{c_{\phi^{\prime}}}{2}E_{\bar{B}T}$
@@
$5^{\prime*}$ $\overline{B}{}^{0}\to T^{0}_{cs}\eta$ $\frac{c_{\phi^{\prime}}}{\sqrt{2}}(C^{\prime}_{T}+C^{\prime}_{TP})-s_{\phi^{\prime}}C^{\prime}_{PT,1}$
$5^{\prime*}$ $\overline{B}{}^{0}\to T^{0}_{cs}\eta^{\prime}$ $\frac{s_{\phi^{\prime}}}{\sqrt{2}}(C^{\prime}_{T}+C^{\prime}_{TP})+c_{\phi^{\prime}}C^{\prime}_{PT,1}$
@@
$6^{\prime*}$ $B^{-}\to T^{-}_{cs}\eta$ $c_{\phi^{\prime}}(C^{\prime}_{T}+C^{\prime}_{TP})-\sqrt{2}s_{\phi^{\prime}}C^{\prime}_{PT,1}$
$6^{\prime*}$ $B^{-}\to T^{-}_{cs}\eta^{\prime}$ $s_{\phi^{\prime}}(C^{\prime}_{T}+C^{\prime}_{TP})+\sqrt{2}c_{\phi^{\prime}}C^{\prime}_{PT,1}$
@@
$11^{\prime*}$ $\overline{B}{}^{0}\to T^{0}_{c}\overline{K}^{0}$ $-\sqrt{\frac{2}{3}}C^{\prime}_{PT,1}$

### v3

@@ V.2.1 $\overline{B}\to TP$ decay amplitudes
$3^{*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}\pi^{0}$ $\frac{1}{\sqrt{2}}(C_{TP}-C_{PT,1})$ $3^{*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}\eta$ $\frac{c_{\theta}}{\sqrt{2}}(C_{TP}+C_{PT,1})-s_{\theta}C_{T}$
$3^{*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}\eta^{\prime}$ $\frac{s_{\theta}}{\sqrt{2}}(C_{TP}+C_{PT,1})+c_{\theta}C_{T}$ $4^{*}$ $\overline{B}{}^{0}\to T^{0}_{cs}K^{0}$ $C_{PT,1}+E_{\bar{B}T}$
55 $\overline{B}{}^{0}\to T^{+}_{c}\pi^{-}$ $\frac{1}{\sqrt{2}}(C_{PT,1}-E_{\bar{B}T})$ 66 $B^{-}\to T^{0}_{c}\pi^{-}$ $\frac{1}{\sqrt{2}}(C_{T}+C_{TP})$
66 $\overline{B}{}^{0}\to T^{0}_{c}\pi^{0}$ $\frac{1}{2}(-C_{T}-C_{TP}+C_{PT,1}-E_{\bar{B}T})$ 66 $\overline{B}{}^{0}\to T^{0}_{c}\eta$ $\frac{c_{\theta}}{2}(C_{T}-C_{TP}-C_{PT,1}+E_{\bar{B}T})+\frac{s_{\theta}}{\sqrt{2}}E_{\bar{B}T}$
66 $\overline{B}{}_{s}^{0}\to T^{0}_{c}K^{0}$ $\frac{1}{\sqrt{2}}(C_{T}+C_{PT,1})$ 66 $\overline{B}{}^{0}\to T^{0}_{c}\eta^{\prime}$ $\frac{s_{\theta}}{2}(C_{T}-C_{TP}-C_{PT,1}+E_{\bar{B}T})-\frac{c_{\theta}}{\sqrt{2}}E_{\bar{B}T}$
77 $\overline{B}{}^{0}\to T^{\prime\prime+}_{c}\pi^{-}$ $\frac{1}{\sqrt{2}}(2T_{P}+C_{PT,1}-E_{\bar{B}T}+2E_{\bar{B}P})$ 88 $B^{-}\to T^{\prime\prime 0}_{c}\pi^{-}$ $\frac{1}{\sqrt{2}}(2T_{P}+C_{T}+C_{TP}+2C_{PT,2})$
88 $\overline{B}{}^{0}\to T^{\prime\prime 0}_{c}\eta$ $\frac{c_{\theta}}{2}(-C_{T}+C_{TP}+C_{PT,1}-2C_{PT,2}$ 88 $\overline{B}{}^{0}\to T^{\prime\prime 0}_{c}\eta^{\prime}$ $\frac{s_{\theta}}{2}(-C_{T}+C_{TP}+C_{PT,1}-2C_{PT,2}$
$E_{\bar{B}T}-2E_{\bar{B}P})+\frac{s_{\theta}}{\sqrt{2}}E_{\bar{B}T}$ $E_{\bar{B}T}-2E_{\bar{B}P})-\frac{c_{\theta}}{\sqrt{2}}E_{\bar{B}T}$
88 $\overline{B}{}^{0}\to T^{\prime\prime 0}_{c}\pi^{0}$ $\frac{1}{2}(C_{T}+C_{TP}-C_{PT,1}+2C_{PT,2}$ 88 $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}K^{0}$ $\frac{1}{\sqrt{2}}(C_{T}-C_{PT,1}+2C_{PT,2})$
$E_{\bar{B}T}-2E_{\bar{B}P})$
99 $\overline{B}{}^{0}\to T^{\prime\prime+}_{c\bar{s}}K^{-}$ $\frac{1}{\sqrt{2}}(C_{TP}+E_{\bar{B}T}-2E_{\bar{B}P})$ 99 $\overline{B}{}_{s}^{0}\to T^{\prime\prime+}_{c\bar{s}}\pi^{-}$ $\frac{1}{\sqrt{2}}(2T_{P}+C_{TP}+C_{PT,1})$
# Mode $A^{\prime}(\overline{B}_{q}\to T_{cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}]}P)$ # Mode $A^{\prime}(\overline{B}_{q}\to T_{cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}]}P)$
$2^{*}$ $\overline{B}{}_{s}^{0}\to T^{+}_{c\bar{s}}K^{-}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{PT,1}-E^{\prime}_{\bar{B}T})$ $3^{*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}\overline{K}^{0}$ $C^{\prime}_{PT,1}+E^{\prime}_{\bar{B}T}$
$4^{*}$ $B^{-}\to T^{0}_{cs}\pi^{-}$ $C^{\prime}_{T}+C^{\prime}_{TP}$ $4^{*}$ $\overline{B}{}^{0}\to T^{0}_{cs}\pi^{0}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{T}+C^{\prime}_{TP})$
$4^{*}$ $\overline{B}{}^{0}\to T^{0}_{cs}\eta$ $\frac{c_{\theta}}{\sqrt{2}}(C^{\prime}_{T}-C^{\prime}_{TP})+s_{\theta}C^{\prime}_{PT,1}$ $4^{*}$ $\overline{B}{}^{0}\to T^{0}_{cs}\eta^{\prime}$ $\frac{s_{\theta}}{\sqrt{2}}(C^{\prime}_{T}-C^{\prime}_{TP})-c_{\theta}C^{\prime}_{PT,1}$
$4^{*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{cs}K^{0}$ $C^{\prime}_{T}+E^{\prime}_{\bar{B}T}$ 55 $\overline{B}{}^{0}\to T^{+}_{c}K^{-}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{TP}-C^{\prime}_{PT,1})$
55 $\overline{B}{}_{s}^{0}\to T^{+}_{c}\pi^{-}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{TP}-E^{\prime}_{\bar{B}T})$ 66 $B^{-}\to T^{0}_{c}K^{-}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{T}+C^{\prime}_{TP})$
66 $\overline{B}{}^{0}\to T^{0}_{c}\overline{K}^{0}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{T}+C^{\prime}_{PT,1})$ 66 $\overline{B}{}_{s}^{0}\to T^{0}_{c}\pi^{0}$ $\frac{1}{2}(C^{\prime}_{TP}-E^{\prime}_{\bar{B}T})$
66 $\overline{B}{}_{s}^{0}\to T^{0}_{c}\eta$ $\frac{c_{\theta}}{2}(C^{\prime}_{TP}+E^{\prime}_{\bar{B}T})$ 66 $\overline{B}{}_{s}^{0}\to T^{0}_{c}\eta^{\prime}$ $\frac{s_{\theta}}{2}(C^{\prime}_{TP}+E^{\prime}_{\bar{B}T})$
$\frac{s_{\theta}}{\sqrt{2}}(C^{\prime}_{T}-C^{\prime}_{PT,1}+E^{\prime}_{\bar{B}T})$ $\frac{c_{\theta}}{\sqrt{2}}(C^{\prime}_{T}-C^{\prime}_{PT,1}+E^{\prime}_{\bar{B}T})$
77 $\overline{B}{}^{0}\to T^{\prime\prime+}_{c}K^{-}$ $\frac{1}{\sqrt{2}}(2T^{\prime}_{P}+C^{\prime}_{TP}+C^{\prime}_{PT,1})$ 77 $\overline{B}_{s}{}^{0}\to T^{\prime\prime+}_{c}\pi^{-}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{TP}+E^{\prime}_{\bar{B}T}-2E^{\prime}_{\bar{B}P})$
88 $B^{-}\to T^{\prime\prime 0}_{c}K^{-}$ $\frac{1}{\sqrt{2}}(2T^{\prime}_{P}+C^{\prime}_{T}+C^{\prime}_{TP}+2C^{\prime}_{PT,2})$ 88 $\overline{B}{}^{0}\to T^{\prime\prime 0}_{c}\overline{K}^{0}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{T}-C^{\prime}_{PT,1}+2C^{\prime}_{PT,2})$
88 $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}\eta$ $\frac{c_{\theta}}{2}(C^{\prime}_{T}-E^{\prime}_{\bar{B}T}-2E^{\prime}_{\bar{B}P})$ 88 $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}\eta^{\prime}$ $\frac{s_{\theta}}{2}(C^{\prime}_{T}-E^{\prime}_{\bar{B}T}-2E^{\prime}_{\bar{B}P})$
$\frac{s_{\theta}}{\sqrt{2}}(C^{\prime}_{T}-C^{\prime}_{PT,1}+2C^{\prime}_{PT,2}+E^{\prime}_{\bar{B}T})$ $\frac{c_{\theta}}{\sqrt{2}}(C^{\prime}_{T}-C^{\prime}_{PT,1}+2C^{\prime}_{PT,2}+E^{\prime}_{\bar{B}T})$
88 $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}\pi^{0}$ $\frac{1}{2}(C^{\prime}_{TP}+E^{\prime}_{\bar{B}T}-2E^{\prime}_{\bar{B}P})$ 99 $\overline{B}{}_{s}^{0}\to T^{\prime\prime+}_{c\bar{s}}K^{-}$ $\frac{1}{\sqrt{2}}(2T^{\prime}_{P}+C^{\prime}_{PT,1}-E^{\prime}_{\bar{B}T}+2E^{\prime}_{\bar{B}P})$
@@
$2^{\prime*}$ $\overline{B}{}^{0}\to T^{+}_{c\bar{s}}K^{-}$ $\frac{1}{\sqrt{2}}(C_{TP}-E_{\bar{B}T})$
$2^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{+}_{c\bar{s}}\pi^{-}$ $\frac{1}{\sqrt{2}}(C_{TP}-C_{PT,1})$
@@
$3^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}\eta$ $s_{\theta}C_{T}+\frac{c_{\theta}}{\sqrt{2}}(C_{PT,1}+C_{TP})$
$3^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}\eta^{\prime}$ $c_{\theta}C_{T}+\frac{s_{\theta}}{\sqrt{2}}(C_{PT,1}+C_{TP})$
@@
$11^{\prime*}$ $B^{-}\to T^{0}_{c}\pi^{-}$ $\sqrt{\frac{2}{3}}(C_{T}+C_{TP}-C_{PT,1})$
@@
$11^{\prime*}$ $\overline{B}{}^{0}\to T^{0}_{c}\eta$ $\frac{c_{\theta}}{\sqrt{3}}(C_{T}+C_{TP}+C_{PT,1})$
$11^{\prime*}$ $\overline{B}{}^{0}\to T^{0}_{c}\eta^{\prime}$ $\frac{s_{\theta}}{\sqrt{3}}(C_{T}+C_{TP}+C_{PT,1})$
$11^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c}K^{0}$ $\sqrt{\frac{2}{3}}C_{T}$
@@
$12^{\prime*}$ $B^{-}\to T^{-}_{c}\eta$ $c_{\theta}(C_{T}+C_{TP}+C_{PT,1})$
$12^{\prime*}$ $B^{-}\to T^{-}_{c}\eta^{\prime}$ $s_{\theta}(C_{T}+C_{TP}+C_{PT,1})$
@@
$14^{\prime}$ $\overline{B}{}^{0}\to T^{\prime 0}_{c}\pi^{0}$ $\frac{1}{2\sqrt{6}}(C_{T}-C_{TP}+C_{PT,1}-E_{\bar{B}T})$
$14^{\prime}$ $\overline{B}{}^{0}\to T^{\prime 0}_{c}\eta$ $\frac{c_{\theta}}{2\sqrt{6}}(C_{T}+C_{TP}+C_{PT,1}+3E_{\bar{B}T})+\frac{\sqrt{3}s_{\theta}}{2}E_{\bar{B}T}$
$14^{\prime}$ $\overline{B}{}^{0}\to T^{\prime 0}_{c}\eta^{\prime}$ $\frac{s_{\theta}}{2\sqrt{6}}(C_{T}+C_{TP}+C_{PT,1}+3E_{\bar{B}T})-\frac{\sqrt{3}c_{\theta}}{2}E_{\bar{B}T}$
@@
$17^{\prime}$ $\overline{B}{}^{0}\to T^{\prime\prime 0}_{c}\eta$ $\frac{c_{\theta}}{2\sqrt{2}}(C_{T}+C_{TP}+C_{PT,1}+4C_{PT,2}+3E_{\bar{B}T}+4E_{\bar{B}P})-\frac{s_{\theta}}{2}E_{\bar{B}T}$
$17^{\prime}$ $\overline{B}{}^{0}\to T^{\prime\prime 0}_{c}\eta^{\prime}$ $\frac{s_{\theta}}{2\sqrt{2}}(C_{T}+C_{TP}+C_{PT,1}+4C_{PT,2}+3E_{\bar{B}T}+4E_{\bar{B}P})+\frac{c_{\theta}}{2}E_{\bar{B}T}$
@@
$5^{\prime*}$ $\overline{B}{}^{0}\to T^{0}_{cs}\eta$ $\frac{c_{\theta}}{\sqrt{2}}(C^{\prime}_{T}+C^{\prime}_{TP})-s_{\theta}C^{\prime}_{PT,1}$
$5^{\prime*}$ $\overline{B}{}^{0}\to T^{0}_{cs}\eta^{\prime}$ $\frac{s_{\theta}}{\sqrt{2}}(C^{\prime}_{T}+C^{\prime}_{TP})+c_{\theta}C^{\prime}_{PT,1}$
@@
$6^{\prime*}$ $B^{-}\to T^{-}_{cs}\eta$ $c_{\theta}(C^{\prime}_{T}+C^{\prime}_{TP})-\sqrt{2}s_{\theta}C^{\prime}_{PT,1}$
$6^{\prime*}$ $B^{-}\to T^{-}_{cs}\eta^{\prime}$ $s_{\theta}(C^{\prime}_{T}+C^{\prime}_{TP})+\sqrt{2}c_{\theta}C^{\prime}_{PT,1}$
@@
$11^{\prime*}$ $\overline{B}{}^{0}\to T^{0}_{c}\overline{K}^{0}$ $\sqrt{\frac{2}{3}}C^{\prime}_{PT,1}$